suppose the random variable x follows a normal distibution with mean 53 and standard deviation 10

a) P(X<43)

b) P(X<60)

c) P(48<X<68)
Answer #1

Given that, mean (μ) = 53 and standard deviation (\sigma) = 10

a) We want to find, P(X < 43)

P(X < 43) = P(\frac {X -\mu}{\sigma} < \frac { 43-53}{10} ) = P(Z < -1) = 0.1587

=> P(X < 43) = 0.1587

b) We want to find, P(X < 60)

P(X < 60) = P(\frac {X -\mu}{\sigma} < \frac { 60-53}{10} ) = P(Z < 0.7) = 0.7580

=> P(X < 60) = 0.7580

c) We want to find, P(48 < X < 68)

\\ P(48 < X < 68) \\\\ = P(\frac { 48-53}{10} < \frac {X -\mu}{\sigma} < \frac { 68-53}{10} ) \\\\ = P(-0.5 < Z < 1.5) \\\\ = P(Z < 1.5) - P(Z < -0.5) \\\\ = 0.9332 - 0.3085 \\\\ = 0.6247

=> P(48 < X < 68) = 0.6247

​​​​​​

​​​​​​​​​